# Scalable Kernel Logistic Regression with Nyström Approximation: Theoretical Analysis and Application to Discrete Choice Modelling

### v1

@@ 4.2 Column selection methods
The previous theoretical result indicates that once the $\mathbf{C}$ matrix is given, the Nyström approximation is optimal. It is for this reason that the effort is expended on finding optimal strategies to select the sketching matrix $\mathbf{P}$. The optimal selection for matrix $\mathbf{P}$ involves computing the $C$ eigenvectors corresponding to the largest eigenvalues $\sigma_{1},\cdots,\sigma_{C}$. Regrettably, achieving high-precision computation of these eigenvectors is computationally demanding. However, a viable alternative is to resort to column selection methods.
Columns selection obtains $\mathbf{C}\in\mathbb{R}^{N\times C}$ using a randomly sampled and rescaled subset of the columns of $\mathbf{K}\in\mathbb{R}^{N\times N}$. The columns of the sketch are selected identically and independently as follows: Let $q_{1},\cdots q_{N}\in(0,1)$, with $q_{1}+\cdots+q_{N}=1$, be the sampling probabilities associated with the columns of $\mathbf{K}$, each column of $\mathbf{C}$ is randomly sampled from the columns of $\mathbf{K}$ according to the sampling probabilities and rescaled by $\frac{1}{\sqrt{Cp_{i}}}$. Three column selection techniques are detailed below.
@@ 4.3 Nyström KLR error bounds
The proof of this theorem is given in the Appendix B. The norm $\left\|\left[\mathbf{K}^{2}-\widehat{\mathbf{K}^{2}}\right]^{1/2}\right\|_{p}$ depends on spectral decay of $\mathbf{K}$ and Theorem 4.1 quantifies the performance of the Nyström KLR method based on this value.
@@ 5.1 Calculating the gradient and loss function in Nyström-based KLR
$\nabla\mathcal{L}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=\frac{1}{N}\mathbf{K}\left[N\lambda\boldsymbol{\alpha}_{\cdot\cdot}-\mathbf{p}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})+\boldsymbol{y}_{\cdot\cdot}\right]\in\mathbb{R}^{N\times I}.$ (26)
Generally, when the sample size $N$ reaches tens of thousands, the computational time becomes prohibitively expensive, posing a significant challenge in handling such large datasets with KLR models. Efficiently computing the matrix of latent functions $\mathbf{f}_{\cdot\cdot}$ is crucial for evaluating the objective function and its gradient during the training of KLR. This is the motivation behind the use of the Nyström approximation, and below is given an outline of its application to the calculation of the gradient and loss function.
@@
$\widehat{\mathbf{p}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=$ $\left[\exp(\widehat{\mathbf{f}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot}))\mathbf{1}\right]^{\top}_{\circ-1}\exp(\widehat{\mathbf{f}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})).$ (28)
$\nabla\mathcal{L}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=$ $\frac{1}{N}\,\mathbf{C}\left[\mathbf{W}^{\dagger}\left[\mathbf{C}^{\top}\left[N\lambda\boldsymbol{\alpha}_{\cdot\cdot}-\widehat{\mathbf{p}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})+\boldsymbol{y}_{\cdot\cdot}\right]\right]\right].$ (29)
@@ 6.1.2 Proposed classifiers
    Support Vector Machine (SVM). This is a binary classifier that employs a kernel to transform the data into a high-dimension space, aiming to identify the optimal linear decision surface (hyperplane) that separates data into two classes. This paper incorporates two variants of this model: one with a linear kernel, a model frequently included in prior studies comparing MNL and ML methods, and another SVM equipped with a radial basis function kernel. The latter model presents estimation problems due to the large size of the datasets used. This is addressed by the use of a strategy similar to that used for KLR, applying the Nÿstrom method. This is done using the same configuration as in the base Nÿstrom KLR model for each dataset. A one-vs-one strategy is applied to extend SVM to multi-class classification. The one-vs-one classification strategy trains a binary SVM classifier for each unique pair of classes and predicts the class that receives the most votes. Note that SVM does not directly provide probability estimates for each class. Hence, the probabilities were approximated using Platt’s method, which consists of training the SVM using $5-$fold CV, and subsequently applying a logistic regression to the SVM scores.
@@ 6.1.3 Model validation
The datasets used in this study were partitioned into training and testing subsets. For the NTS dataset, a random split was made, with $70\%$ of the data used for training and $30\%$ for testing. In line with the methodology employed by Hillel et al. (2018), the LPMC dataset was divided into a training set comprising the first two years of data (approximately $70\%$ of the dataset) and a test set comprising the last year of data. Using the training set as reference, all variables in both training and test sets were were normalised, centering them to zero mean and unit standard deviation. It should be noted that no measures were taken to address the class imbalance in the datasets.
@@ 6.2 Comparison of Nyström methods for KLR on large datasets
The primary objective of the initial experiment was to evaluate the GMPCA, DCA, and the computational cost of various implementations of KLR, using the column selection methods presented in Section 4.2. The parameters of these models were estimated using the L-BFGS-B algorithm. Specifically, the comparison focuses on determining the number of landmark points $C$ required in each Nyström approximation to achieve satisfactory results.
Before exploring the performance of various column selection methods in the KLR, the quality of the best Nyström approximation of rank $C$ to the kernel matrix $\mathbf{K}$ is considered. Figure 1 illustrates the decay of the spectrum of matrix $\mathbf{K}$ across both datasets used in this study. The optimal Nyström approximation of rank $C$ is $\widehat{\mathbf{K}}_{C}=\sum_{j=1}^{C}\sigma_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}$, where the spectral norm of the residual $\|\mathbf{K}-\widehat{\mathbf{K}}_{C}\|_{\infty}=\sigma_{C+1}$. Figure 1 displays a similar trend for both problems, despite the datasets having entirely distinct characteristics. It is observed that, by selecting more than $500$ landmarks, Nyström matrix approximations of the order $10^{-1}$ are achievable. For the NTS problem, more landmark points are required to achieve the same level of accuracy as for the LPMC.
@@
In both datasets, as the number of landmark points in the Nyström method increases, all the methods eventually converge to a similar level of accuracy, equal to the optimal rank $C$ Nyström approximation $\widehat{\mathbf{K}}_{C}$, approaching the performance that would be achieved by directly applying KLR without the Nyström approximation. However, it is important to highlight that using KLR directly, without Nyström, on these datasets would not be feasible due to the size of the kernel matrix. To illustrate this, storing the training matrices ($70\%$ of the observations) of the LPMC and NTS datasets would require approximately $22$ GB and $194$ GB of RAM memory, respectively. Hence, the adoption of the techniques described here becomes indispensable for handling such large datasets.
@@
The analysis of the results suggests that the $k$-means Nyström KLR method exhibits faster convergence on both datasets, especially when the number of landmark points is small. This can be evidenced by the lower values of the GMPCA and DCA indices when using less than $100$ landmark points. Additionally, the results of the NTS dataset indicate that, as the dataset size increases, this method becomes more effective when working with few landmark points. However, as the number of landmark points grows larger, there is little discernible difference between the various techniques, except for a slight overall improvement in GMPCA and DCA. Nevertheless, it should be noted that the estimation time of the KLR model and the size of the kernel matrix will significantly increase when using a large number of landmark points.
@@ 6.3 Comparison of the Nyström KLR method with other ML algorithms on large datasets
The results of this experiment are set out in Table 1. As seen in Figure 2 for the LPMC dataset, the Nyström KLR with $500$ landmark points returns asymptotic values for the DCA and GMPCA, so there is hardly any difference in performance between the models. Hence, $500$ landmark points were used for the Nyström KLR models. For the NTS dataset, a much larger dataset, Figure 3 shows that for $500$ landmark points the DCA and GMPCA indices still have a small positive slope, therefore $1,000$ landmark points were chosen for the Nyström methods. If the number of landmark points is reduced, it can be seen that the performance of the $k$-means Nyström method would decrease slightly, while the other Nyström methods would experience a more severe decrease.
@@
In terms of the computational time required to estimate the Nyström KLR models, although they are still significant, they are much lower than those of the linear SVM model, especially for very large datasets such as NTS, and similar to those of the SVM with Nÿstrom. Furthermore, the estimation time of all Nyström KLR models, except the DAC ridge-leverage, is lower than the time required by the MNL models, commonly used by practitioners in this field, for such large datasets.
@@
Regarding the KLR models, the best KLR implementation for each dataset is marked with an asterisk ‘*’. It can be observed that for the LPMC dataset, which is the smallest one, the results of all the Nyström KLR models are quite close to those of the top-performing ML algorithms, which suggests they are promising models. However, by employing sampling strategies, the Nystrom-based KLR models demonstrate improved performance compared to the base MNL model. It is noteworthy that, in the NTS dataset, the Nyström KLR methods yield lower scores in terms of DCA. Nevertheless, they still provide better approximations to the probability of each alternative, as measured through the GMPCA, compared to the MNL models.
@@
Another possible explanation for the performance difference is that the LPMC dataset focuses on discrete choice modelling, whereas the NTS dataset is more oriented towards ML applications, and not all variables are equally important. For instance, the NTS dataset contains a significant number of environmental variables, which might have a reduced impact on the classification process. This issue was highlighted in Martín-Baos et al. (2023), where it was calculated the SHAP values for each variable in both datasets, which allowed the variations in variable importance to be analysed. In the LPMC dataset, numerous variables influence the decision maker, while in the NTS dataset, only a few variables, primarily the ‘distance’, have a notable impact. This highlights the sensitivity of the KLR method to non-significant features, suggesting the importance of conducting an attribute analysis prior to applying these techniques. The poor performance of the Nyström SVM on this dataset may also be attributed to these reasons, as well as to the need for a greater number of landmark points for this model.
@@
6.4 Comparative of optimisation techniques for Nyström KLR
After overcoming the challenge of storing large kernel matrices, the subsequent step involves determining the most efficient optimisation algorithm to estimate the vector of parameters in KLR. The estimation process involves solving the optimisation problem in Equation (7), where the loss function is the negative conditional log-likelihood, leading to a procedure called minimising RNLL or, equivalently, maximising PMLE. Typically, the algorithms used for solving this PMLE problem rely on the gradient of the objective function.
From a theoretical perspective, training MNL or KLR models using PMLE is equivalent, as both require estimating functions that are linear in their parameters. However, the differences appear from a computational point of view due to the number of parameters in the models. In the case of MNL (in the context of RUM), the number of parameters increases on the order of ${\cal O}(M\cdot I)$, where $M$ is the dimension of the original input space, i.e., $\mathbf{X}\subset\mathbb{R}^{M}$, and $I$ is the number of alternatives in the choice set. On the other hand, KLR grows more abruptly with ${\cal O}(N\cdot I)$, where $N$ is the number of observations. Consequently, training KLR models poses a computational challenge, as it requires the estimation of a large number of coefficients.

### v3

@@ 4.2 Column selection methods
The previous theoretical result indicates that once the $\mathbf{C}$ matrix is given, the Nyström approximation is optimal. It is for this reason that the effort is expended on finding optimal strategies to select the sketching matrix $\mathbf{P}$. The sketching matrix for best $C-$rank approximation is given by ${\mathbf{P}}=(\mathbf{U}_{\cdot 1},\ldots,\mathbf{U}_{\cdot C}){\rm´diag}(\frac{1}{\sigma_{1}},\ldots,\frac{1}{\sigma_{C}})$ where $\mathbf{U}_{\cdot j}$ are the first elements of the orthonormal basis used in the SVD decomposition given in Equation (9), assuming that ${\rm rank(\mathbf{C})}=C$. This involves computing the $C$ eigenvectors corresponding to the largest eigenvalues $\sigma_{1},\cdots,\sigma_{C}$. Regrettably, achieving high-precision calculation of these eigenvectors is computationally demanding. However, a viable alternative is to resort to column selection methods.
Columns selection obtains $\mathbf{C}\in\mathbb{R}^{N\times C}$ using a randomly sampled and rescaled subset of the columns of $\mathbf{K}\in\mathbb{R}^{N\times N}$. The columns of the sketch are selected identically and independently as follows: Let $q_{1},\cdots q_{N}\in(0,1)$, with $q_{1}+\cdots+q_{N}=1$, be the sampling probabilities associated with the columns of $\mathbf{K}$, each column of $\mathbf{C}$ is randomly sampled from the columns of $\mathbf{K}$ according to the sampling probabilities and rescaled by $\frac{1}{\sqrt{Cq_{i}}}$, where $q_{i}$ is the sampling probability of column $i$. Three column selection techniques are detailed below.
@@ 4.3 Nyström KLR error bounds
The proof of this theorem is given in the B. The norm $\left\|\left[\mathbf{K}^{2}-\widehat{\mathbf{K}^{2}}\right]^{1/2}\right\|_{p}$ depends on spectral decay of $\mathbf{K}$ and Theorem 4.1 quantifies the performance of the Nyström KLR method based on this value.
@@ 5.1 Calculating the gradient and loss function in Nyström-based KLR
$\nabla\mathcal{L}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=\frac{1}{N}\mathbf{K}\left[N\lambda\boldsymbol{\alpha}_{\cdot\cdot}+\mathbf{p}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})-\boldsymbol{y}_{\cdot\cdot}\right]\in\mathbb{R}^{N\times I}.$ (26)
Generally, when the sample size $N$ reaches tens of thousands, the computational time becomes prohibitively expensive, posing a significant challenge in handling such large datasets with KLR models. Efficiently computing the matrix of latent functions $\mathbf{f}_{\cdot\cdot}$ is crucial for obtaining $\mathbf{p}_{\cdot\cdot}$, as stated in Equation (6), and therefore evaluating the objective function and its gradient during the training of KLR. This is the motivation behind the use of the Nyström approximation, and below is given an outline of its application to the calculation of the gradient and loss function.
@@
$\widehat{\mathbf{p}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=$ $\left[\left(\exp(\widehat{\mathbf{f}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot}))\mathbf{1}\right)_{\circ-1}\right]^{\top}\exp(\widehat{\mathbf{f}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})).$ (28)
$\nabla\mathcal{L}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})=$ $\frac{1}{N}\,\mathbf{C}\left[\mathbf{W}^{\dagger}\left[\mathbf{C}^{\top}\left[N\lambda\boldsymbol{\alpha}_{\cdot\cdot}+\widehat{\mathbf{p}}_{\cdot\cdot}(\boldsymbol{\alpha}_{\cdot\cdot})-\boldsymbol{y}_{\cdot\cdot}\right]\right]\right].$ (29)
@@ 6.1.2 Proposed classifiers
    Support Vector Machine (SVM). This is a binary classifier that employs a kernel to transform the data into a high-dimension space, aiming to identify the optimal linear decision surface (hyperplane) that separates data into two classes. This paper incorporates two variants of this model: one with a linear kernel, a model frequently included in prior studies comparing MNL and ML methods, and another SVM equipped with a radial basis function kernel. The latter model presents estimation problems due to the large size of the datasets used. This is addressed by the use of a strategy similar to that used for KLR, applying the Nyström method. This is done using the same configuration as in the base Nyström KLR model for each dataset. A one-vs-one strategy is applied to extend SVM to multi-class classification. The one-vs-one classification strategy trains a binary SVM classifier for each unique pair of classes and predicts the class that receives the most votes. Note that SVM does not directly provide probability estimates for each class. Hence, the probabilities were approximated using Platt’s method, which consists of training the SVM using $5-$fold CV, and subsequently applying a logistic regression to the SVM scores.
@@ 6.1.3 Model validation
The datasets used in this study were partitioned into training and testing subsets. For the NTS dataset, a random split was made, with $70\%$ of the data used for training and $30\%$ for testing. In line with the methodology employed by Hillel et al. (2018), the LPMC dataset was divided into a training set comprising the first two years of data (approximately $70\%$ of the dataset) and a test set comprising the last year of data. Using the training set as reference, all variables in both training and test sets were normalised, centering them to zero mean and unit standard deviation. It should be noted that no measures were taken to address the class imbalance in the datasets.
@@ 6.2 Comparison of Nyström methods for KLR on large datasets
Theorem 4.1 shows that the convergence of the parameters of the Nyström KLR to the parameters of the KLR depends on the norm $\|(\mathbf{K}^{2}-\widehat{\mathbf{K}}^{2})^{1/2}\|_{p}$. More specifically, if ${\cal E}(C)$ is the Schatten $p-$norm of the error in the parameters of the Nyström KLR, which depends on the number of landmark points $C$, it is proved that ${\cal E}(C)=O(\|(\mathbf{K}^{2}-\widehat{\mathbf{K}}^{2})^{1/2}\|_{p})$.
The behaviour of this bound will be analysed numerically for two transport mode choice problems. The theorem is satisfied for any Nyström approximation, in particular for the best $C-$rank approximation, which is given by the expression $\widehat{\mathbf{K}}_{C}=\sum_{j=1}^{C}\sigma_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}$, obtaining:
$(\mathbf{K}^{2}-\widehat{\mathbf{K}}_{C}^{2})^{1/2}$ $=\left(\sum_{j=1}^{R}\sigma^{2}_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}-\sum_{j=1}^{C}\sigma^{2}_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}\right)^{1/2}=\left(\sum_{j=C+1}^{R}\sigma^{2}_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}\right)^{1/2}$ (38)
$=\sum_{j=C+1}^{R}\sigma_{j}\mathbf{U}_{\cdot j}\mathbf{U}_{\cdot j}^{\top}\Rightarrow\|(\mathbf{K}^{2}-\widehat{\mathbf{K}}_{C}^{2})^{1/2}\|_{p}=\left(\sum_{j=C+1}^{R}\sigma^{p}_{j}\right)^{1/p}.$ (39)
Considering the spectral norm ($p=\infty$), it follows that $\|(\mathbf{K}^{2}-\widehat{\mathbf{K}}_{C}^{2})^{1/2}\|_{\infty}=\sigma_{C+1}$. This leads to the fact that the error bound, as a function of the number of landmark points $C$, has an expression of the form ${\cal E}(C)\leq\delta\sigma_{C+1}$.
Figure 1 shows the decay of the spectrum of $\mathbf{K}$, $\sigma_{C+1}$ versus $C$, in the two datasets. Therefore, the error bounds of the parameters will be proportional to the functions shown in Figure 1. Fitting the regression model $\sigma_{C+1}=\dfrac{a}{C^{b}}$, the fitted model $\sigma_{C+1}=\dfrac{e^{6.58}}{C^{1.36}}$ is obtained for the LPMC problem, and $\sigma_{C+1}=\dfrac{e^{6.51}}{C^{1.30}}$ for the NTS dataset. These results indicate a rapid spectral decay ($b>1$) in these application examples, suggesting that a large number of landmarks is not required for an effective Nyström approximation. Additionally, it is observed that $\|(\mathbf{K}^{2}-\widehat{\mathbf{K}}_{C}^{2})^{1/2}\|_{p}=\|\mathbf{K}-\widehat{\mathbf{K}}_{C}\|_{p}$, and thus by selecting more than $500$ landmarks, the error of Nyström matrix approximations with respect to the kernel matrix $\mathbf{K}$ is of the order of $\sigma_{500}\approx 10^{-1}$.
@@
Next, GMPCA, DCA, and the computational cost of various implementations of KLR, are evaluated, using the column selection methods presented in Section 4.2. For each implementation of KLR, a Radial Basis Function (RBF) kernel was selected, and the parameters of these models were estimated using the L-BFGS-B algorithm. Specifically, the comparison focuses on determining the number of landmark points $C$ required in each Nyström approximation to achieve satisfactory results. The question that arises is whether the new metrics exhibit the same pattern as that registered in this experiment. Intuition indicates that they should be similar and that therefore these metrics should be seen to give small improvements as they go beyond $300$ landmarks.
@@
In both datasets, as the number of landmark points in the Nyström method increases, all the methods eventually converge to a similar level of accuracy. When $C$ grows, the standard Nyström KLR converges to the best rank $C-$approximation $\widehat{\mathbf{K}}_{C}$ (Kumar et al. (2012), Theorem 2) and in turn converges to the kernel matrix $\mathbf{K}$ (Sun et al. (2015), Theorem 6). Therefore, since Figures 2 and 3 shows that all Nyström schemes (including random sampling) converge to the same level of accuracy, it is this performance that would be achieved by directly applying KLR. However, it is important to highlight that using KLR directly, without Nyström, on these datasets would not be feasible due to the size of the kernel matrix. To illustrate this, storing the training matrices ($70\%$ of the observations) of the LPMC and NTS datasets would require approximately $22$ GB and $194$ GB of RAM memory, respectively. Hence, the adoption of the techniques described here becomes indispensable for handling such large datasets.
@@
The analysis of the results suggests that the $k$-means Nyström KLR method exhibits faster convergence on both datasets, especially when the number of landmark points is small. This can be evidenced by the higher values of the GMPCA and DCA indices when using less than $100$ landmark points. Additionally, the results of the NTS dataset indicate that, even with a larger dataset, the $k$-means Nyström method achieves convergence using only a few landmark points. However, as the number of landmark points grows larger, there is little discernible difference between the various techniques, except for a slight overall improvement in GMPCA and DCA. Nevertheless, it should be noted that the estimation time of the KLR model and the size of the kernel matrix will significantly increase when using a large number of landmark points.
To conclude this experiment, it should be noted that the metrics GMPCA and DCA behave similarly to the error bound of the parameters ${\cal\varepsilon}(C)$. Looking at Figures 2 and 3, which employ logarithmic scaling on the X-axis, the performance improvement is negligible above $300$ landmark points.
@@ 6.3 Comparison of the Nyström KLR method with other ML algorithms on large datasets
The results of this experiment are set out in Table 1. For each implementation of KLR, a RBF kernel was employed, and the model parameters were estimated using the L-BFGS-B algorithm. As seen in Figure 2 for the LPMC dataset, the Nyström KLR with $500$ landmark points returns asymptotic values for the DCA and GMPCA, so there is hardly any difference in performance between the models. Hence, $500$ landmark points were used for the Nyström KLR models. For the NTS dataset, a much larger dataset, Figure 3 shows that for $500$ landmark points the DCA and GMPCA indices still have a small positive slope, therefore $1,000$ landmark points were chosen for the Nyström methods. If the number of landmark points is reduced, it can be seen that the performance of the $k$-means Nyström method would decrease slightly, while the other Nyström methods would experience a more severe decrease.
@@
Computational time, also referred to as training time, is an important factor to consider. In terms of the time required to estimate the Nyström KLR models, although it remains substantial, it is significantly lower than that of the linear SVM model, especially for very large datasets like NTS. The time required is also comparable to that of the Nyström SVM model. Moreover, with the exception of the DAC ridge-leverage variant, all Nyström KLR models require less time to estimate than the MNL models, which are commonly used by practitioners for large datasets.
@@
Regarding the KLR models, the best KLR implementation for each dataset is marked with an asterisk ‘*’. It can be observed that for the LPMC dataset, which is the smallest one, the results of all the Nyström KLR models are quite close to those of the top-performing ML algorithms, which suggests they are promising models. However, by employing sampling strategies, the Nyström-based KLR models demonstrate improved performance compared to the base MNL model. It is noteworthy that, in the NTS dataset, the Nyström KLR methods yield lower scores in terms of DCA. Nevertheless, they still provide better approximations to the probability of each alternative, as measured through the GMPCA, compared to the MNL models.
@@
Another possible explanation for the performance difference is that the LPMC dataset focuses on discrete choice modelling, whereas the NTS dataset is more oriented towards ML applications, and not all variables are equally important. For instance, the NTS dataset contains a significant number of environmental variables, which might have a reduced impact on the classification process. This issue was highlighted in Martín-Baos et al. (2023), where it was calculated the SHAP values for each variable in both datasets, which allowed the variations in variable importance to be analysed. In the LPMC dataset, a wide range of variables significantly influences the decision-maker’s choices. Conversely, in the NTS dataset, despite having fewer variables, only a few of them, primarily the ‘distance’, have a notable impact on the classification. The remaining variables are less significant and can impact classifier performance. This accounts for the generally lower effectiveness of all classifiers, particularly the KLR. This highlights the sensitivity of the KLR method to non-significant features, suggesting the importance of conducting an attribute analysis prior to applying these techniques.
@@
Note: For Nyström KLR models, $500$ landmark points were used for the LPMC dataset and $1,000$ for the NTS dataset.
@@
Note: For Nyström KLR models, $500$ landmark points were used for the LPMC dataset and $1,000$ for the NTS dataset.
6.4 Comparison of optimisation techniques for Nyström KLR
After overcoming the challenge of storing large kernel matrices, the subsequent step involves determining the most efficient optimisation algorithm to estimate the vector of parameters in KLR. The estimation process involves solving the optimisation problem in Equation (7), where the loss function is the negative conditional log-likelihood, leading to a procedure called minimising RNLL or, equivalently, maximising PMLE. Typically, the algorithms used for solving this PMLE problem rely on the gradient of the objective function. From a theoretical point of view, training MNL or KLR models with PMLE is fundamentally equivalent, as both are instances of the same Generalised Linear Model (GLM). However, the differences appear from a computational point of view due to the number of parameters in the models. In the case of MNL (in the context of RUM), the number of parameters increases on the order of ${\cal O}(M\cdot I)$, where $M$ is the dimension of the original input space, i.e., $\mathbf{X}\subset\mathbb{R}^{M}$, and $I$ is the number of alternatives in the choice set. On the other hand, KLR grows more abruptly with ${\cal O}(N\cdot I)$, where $N$ is the number of observations. Consequently, training KLR models poses a computational challenge, as it requires the estimation of a large number of coefficients.